What is the standard deviation of return for a stock with annual return of 13%, 9%, -12%, 10% over the last four years.
Answer #1

Solution :- Returns = 13% , 9% , -12% , 10%

Average Return = ( 13% + 9% - 12% + 10% ) / 4 = 20% / 4 = 5%

19 fix B с D E F A 1 Solution :- 2 3 4 Year 1 2 5 6 7 Returns 13.00% 9.00% -12.00% 10.00% D=(R-AR) 8.00% 4.00% -17.00% 5.00%

Variance = D * D / ( n - 1 ) = 3.94% / ( 4 - 1 ) = 1.3133%

Standard Deviation = ( Var )1/2 = ( 1.3133% )1/2 = 11.46%

If there is any doubt please ask in comments

Thank you please rate